OpenStudy (anonymous):

x2 - 8 -------- x + 6 and x =-2

OpenStudy (tkhunny):

Have you considered substitution? x+6 given x = -2 gives (-2) + 6 = 4 You do the numerator. Note: Folks in the chats are interested mostly in talking about other things, rather than helping or studying. Sorry about that. We can't seem to change that behavior.

OpenStudy (anonymous):

This is your new problem once you substitute x for -2 \[\frac{ -2\times2-8 }{ -2+6 }\]

OpenStudy (anonymous):

Do you think you will be able to solve that?

OpenStudy (tkhunny):

I believe that would be \((-2)^{2} - 8\)

OpenStudy (anonymous):

Maybe, the exponent sign was not placed, but most likely your right.

OpenStudy (anonymous):

If yes, this is the problem \[\frac{ -2^2-8 }{ -2+6 }\]

OpenStudy (tkhunny):

Just as a side note, if you are programming computers, the notation "-2^2" can be a little ambiguous. Your specific language should tell you what it means. In some, the "Unary Minus" is very high precedence and you get (-2)^2. In others, evaluation may be strictly right-to-left and you get -(2^2). In my opinion, this situation is important to clarify with parentheses. Make sure whoever is reading your work understands what you mean by -2^2. End of Side Note.